Circular No. 7402

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 2000bk IN NGC 4520 Mark Armstrong, Rolvenden, Kent, England, reports his discovery of

an apparent supernova (mag 17) visible on four CCD images taken between Apr. 11.942 and 12.090 UT. SN 2000bk is located at α = 12^{h}33^{m}53^{s}.94, δ = −7^{o}22′42′′.9 (equinox 2000.0; mean of measurements from four frames), which is 61′′ east and 10′′ south of the nucleus of NGC 4520. Earlier images taken by Armstrong on Jan. 10 and Mar. 4–5 show nothing at this position (limiting mag 18), neither is anything visible on the IIIa-J SERC J survey image taken on 1983 Apr. 15 (limiting mag 23) or on the IIIa-F Palomar Sky Survey plate taken on 1993 Feb. 23 (limiting mag 22).

W. Li, University of California at Berkeley, reports that a prediscovery image of SN 2000bk was obtained with the 0.8-m Katzman Automatic Imaging Telescope on Apr. 7.3 UT, showing the new object at mag ∼ 17.0.

SMC X-2 R. Corbet, Goddard Space Flight Center (GSFC) and Universities

Space Research Association; and F. E. Marshall, GSFC, report on behalf of the RXTE teams at GSFC and the Massachusetts Institute of Technology: “We have discovered pulsed x-ray emission from a position consistent with the location of SMC X-2. RXTE ASM observations indicate activity since the middle of January with a mean flux of ∼ 10 mCrab. Slews with the PCA over SMC X-2 on Apr. 9 and 12 show a source at α = 0^{h}54^{m}53^{s}, δ = −73^{o}38′, with a 90-percent-confidence error circle of 3′ radius and a 2–10-keV flux of ∼ 3 mCrab. A pointed PCA observation on Apr. 12 shows pulsations at a period of 2.374 ± 0.007 s with a pulsed fraction of ∼ 30 percent. Imaging x-ray observations, and observations of the suggested optical counterpart (Murdin et al. 1979, MNRAS 186, 43P), are encouraged.”

UX ANTLIAE P. Williams, Heathcote, New South Wales, reports that this R CrB-

type star appears to be fading from its normal visual mag 12.0, as indicated by his own visual magnitude estimates: Mar. 1.475 UT, 12.0; 3.450, 12.0; 31.473, 12.1; Apr. 11.434, 12.5; 13.534, 12.9. Williams adds that this is the first recorded fade of UX Ant since that reported in June/July 1990 (cf. IAUC 5058, 5089).

2000 April 14  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green